OpenStudy (anonymous):

y= .25x+1 -x+4y=4 Substitution method. I just don't even know with this.

OpenStudy (anonymous):

substitution is basically solve for one variable use the number what you got for that substitute that number back in the remaining formula to get your answer

OpenStudy (anonymous):

Yeah I know but I just don't understand how to do it in this one for some reason. I just keep getting some weird stuff and i just..ughh.

OpenStudy (anonymous):

i have y=2-1y at the end i just don't know what I was doing or what? haha

OpenStudy (anonymous):

can you show me your steps?

OpenStudy (anonymous):

y=.25x+1 -x+4y=4 -x+4y-4y=4-4y -x=4-4y -x/-1=4/-1 -4y/-1 x=4+4y y=.25(4+4y)+1 y=-1+1y+1 y=1y oh y/1=1y/1 y=y?

OpenStudy (anonymous):

Oh good lord, nevermind. Good god -.-